Kandy is financing a $335,000 mortgage for 30 years at a fixed rate of 7.45% what is the total cost of the principal and interest after 30 years


1
Expert's answer
2021-05-13T13:04:15-0400

Calculating the monthly payment:

"=\frac{r\times PV}{1-(1+r)^{-n}}\\ 
\\ 
=\frac{\frac{0.0745}{12}\times\$335,000}{1-(1+\frac{0.0745}{12})^{-30\times 12}}\\=\$2,330.91"


Where:

Present value of the loan (PV) = $335,000

Monthly interest rate ("r) = \frac{0.0745}{12}"

Number of monthly payments (n) = 30 x 12

 

Thus, the monthly payment is $2,330.91

 

Calculating the total cost of principal and interest after 30 years:

"Total\space borrowing\space costs\\=Monthly\space payments\times Total\space number\space of\space payments\\=\$2,330.91\times(30\times12)\\=\$839,127.50"

Thus, the total borrowing cost is $839,127.50.